5.0g of Toluene(C7 H8 ). + 225 g of benzene (C6 H6 ). Density of solution is 0.876 g/mL. Calculate Molarity
1
Expert's answer
2017-11-13T09:16:07-0500
V = m / p = (5 g + 225 g) / 0.876 (g/mL) = 262.55 mL
n (C7H8) = m / M = 5 g / 92 (g/mol) = 0.054 mol
n (C6H6) = m / M = 225 g / 78 (g/mol) = 2.884 mol
C = n / V = (0.054 mol + 2.884 mol) / 0.26255 L = 11.19 mol/L
Answer:
C = 11.19 mol/L